# Self-intersection Number of Negative Curves on Fermat Surfaces

### v2

@@ 3.1. Strategy
$=$ $m^{2}kC^{2}+m^{2}\sum_{i\neq j}C_{i}\cdot C_{j},$
where $C_{i}\cdot C_{j}$ denotes the intersection number on $X_{d}$.
To study the Bounded Negativity Conjecture on $X_{d}$, we need to obtain a lower bound for $C^{2}$. From the formula above, we are led to give an upper bound for the sum
$\sum_{i\neq j}C_{i}\cdot C_{j}.$
The result will be a function of $e$ where $d$ is viewed as a constant, and so is the difference
@@
If this function is bounded below for $e\geq 1$, then the Bounded Negativity Conjecture for $X_{d}$ will follow.
The main part of this paper is to calculate the intersection numbers using local computations. The results are summarized in Proposition 4.4, and the lengthy proof will be presented in Section 6.
Fix an arbitrary intersection point $P=[x_{0},y_{0},z_{0},w_{0}]$ of some $C_{i}$ and $C_{j}$ with $i\neq j$, we will give an explicit formula for the sum
@@
where $I_{P}(C_{i},C_{j})$ denotes the intersection multiplicity (that will be explored in detail in Section 6).
Since we consider all curves $C$ on $X_{d}$, we should view $g$ as an abstract polynomial. From $P\in X_{d}$, we have that $x_{0}^{d}+y_{0}^{d}+z_{0}^{d}+w_{0}^{d}=0$, hence not all $x_{0},y_{0},z_{0}$ are zero. By the symmetry of the coordinates $x,y,z$ in the equation $X_{d}:x^{d}+y^{d}+z^{d}+w^{d}=0$, we may without loss of generality assume that $x_{0}\neq 0$. To decide where the intersection point $P=[x_{0},y_{0},z_{0},w_{0}]$ may situate on $X_{d}$, we propose the following Lemma which will be proved at the end of this section.
@@
Recall that $w_{0}=0$ and we assumed $x_{0}\neq 0$. To calculate the intersection multiplicities at $P$, we introduce affine coordinates
@@
Moreover, since $x_{0}^{d}+y_{0}^{d}+z_{0}^{d}=0$ and $x_{0}\neq 0$, at least one of $y_{0},z_{0}$ does not vanish. We may without loss of generality assume that $z_{0}\neq 0$. Then
@@
Therefore, we can locally express $v$ as an analytic function of $(u,\omega)$ by applying the implicit function theorem to the equation $l(u,v)+\omega^{d}=0$. Then the local equation for $\mathscr{C}$ is $h\left(u,v(u,\omega)\right)=0$. By analyzing this local equation, we can investigate the local branches and compute the local intersection multiplicities. This is the usual way to analyze the curve $\mathscr{C}$.
Our method is, however, first to solve $v$ from the equation $h(u,v)=0$ by using the Puiseux series method (which will be recalled in Section 6), say $v=s(u)$. Then, using the equation $l(u,s(u))+\omega^{d}=0$, we discuss the properties of the curve $\mathscr{C}$ and calculate the intersection multiplicities. To this end, we need that $h$ is $v$-general (the related notions will be discussed in Section 6).
@@
Recall that $h$ is transformed from an irreducible polynomial $g$, thus $h$ is an irreducible polynomial in $u,v$. If $h$ is not $v$-general, then $u$ is a factor of $h$; therefore, by irreducibility, $h$ is a constant multiple of $u$. Back to $g$, we see that $g$ is a linear function $y-ax$ (up to multiplicative factor). Therefore, the curve $\mathscr{C}$ is defined as
@@
If $a^{d}+1\neq 0$, then $\mathscr{C}$ is itself irreducible and hence it must be equal to $C$. Thus, $C^{2}=d$ in this case.
@@
Recall that we have the degree $d$ branched Galois cover $\rho:X_{d}\to\mathbb{P}^{2}$ defined by $[x,y,z,w]\mapsto[x,y,z]$. The ramification locus of $\rho$ is the closed subset of $X_{d}$ defined by the equation $w=0$.
If $Q$ is a point on at least two components of $\mathscr{C}$, then #$\#\left(\rho^{-1}(\rho(Q))\right)<d$, where #$\#A$ means the number of points in a set $A$. It follows that $Q$ lies on the ramification locus of $\rho$. Thus, $w(Q)=0$ as desired.
Furthermore, denote $q=\rho(Q)$, then $\rho^{-1}(q)=\{Q\}$. For all $i\leq k$, the restriction morphism $\rho|_{C_{i}}:C_{i}\to E$ is a finite morphism and thus $\rho^{-1}(q)\cap C_{i}\neq\emptyset$. But $\rho^{-1}(q)$ consists of only one point $Q$, it follows that $Q\in C_{i}$. We are done.
@@ 4.1. Position of intersection points
By definition, $\rho^{-1}(q)=\{Q\}$ and $Q\in\mathscr{C}$. Since each $C_{i}$ is mapped onto $E$, the intersection $\rho^{-1}(q)\cap C_{i}$ is nonempty. Hence $Q\in C_{i}$ as desired.
@@ 4.3. Intersection multiplicities
Recall that we have the following equation
$de^{2}=m^{2}kC^{2}+m^{2}\sum_{i\neq j}C_{i}\cdot C_{j}.$
To give a lower bound for $C^{2}$, we need to estimate the sum on the right-hand side. Moreover, by definition, we have
$\sum_{i\neq j}C_{i}\cdot C_{j}=\sum_{\begin{subarray}{c}i\neq j\\ Q\in C_{i}\cap C_{j}\end{subarray}}I_{Q}(C_{i},C_{j}).$
Hence, it suffices to compute the sum
$\sum_{i\neq j}I_{P}(C_{i},C_{j})$
where $P=[x_{0},y_{0},z_{0},0]$ is chosen above. Our main technical results are summarized as follows.
@@
    $\mu_{P}(\mathscr{C})=(d-1)i_{p}(D,E)+d\mu_{p}(E)-d+1.$
@@
Now we use the above results to study the Bounded Negativity Conjecture for Fermat surfaces. It is clear that we only need to consider negative curves $C$. Then the curve $\mathscr{C}$ necessarily has $k\geq 2$ irreducible components since otherwise $C^{2}$ would be nonnegative. Moreover, the number $k$ divides $d$ since the Galois action for $\rho$ is a cyclic action of order $d$. To ensure that the polynomial $h$ is $v$-general, we only need to assume $C^{2}\neq 2-d$; see the discussions in Section 3. Therefore, combining all of our discussions above, we get the following theorem.
Theorem 4.5.
Let $X_{d}\subset\mathbb{P}^{3}$ be the Fermat surface defined by $x^{d}+y^{d}+z^{d}+w^{d}=0$ for $d\geq 5$, and $C$ a reduced and irreducible curve on $X_{d}$ with $C^{2}<2-d<0$. Then
(1) There exists an irreducible homogeneous polynomial $g(x,y,z)$ which vanishes on $C$.
Denote by $e=\deg g$. Define two curves $E$ and $D$ in $\mathbb{P}^{2}$ by
$E:\ g=0\quad\mathrm{and}\quad D:x^{d}+y^{d}+z^{d}=0.$
(2) There exists $k\geq 2$ and $k|d$ such that
@@
In particular, when $d$ is prime, then $k=d$ and
@@
for some constant $\lambda>0$ which may depend on $d$, where $E$ ranges over all irreducible projective plane curves of degree $e$ and $D$ is the projective plane curve defined by $x^{d}+y^{d}+z^{d}=0$.
$\limsup_{e\to\infty}\sup_{E}\frac{\sum\limits_{p\in D\cap E}\mu_{p}(E)}{e^{2}}<1,$
where again $E$ ranges over all irreducible projective plane curves of degree $e$ and $D$ is as defined above.
@@
By our computations before, it is clear that (a) implies the Bounded Negativity Conjecture. Indeed, if (a) holds, there exists $e_{0}>d+1$ such that
@@
by [12, Cor.7.1.3], where $r_{p}$ is the number of local branches of $E$ at $p$. It follows that
@@
for some $\alpha>0$. The desired result follows by noting that
@@
5. Discussions and Problems
In this section, we investigate some cases in which we are able to estimate the intersection number $C^{2}$ which is equal to
$\frac{d}{k}e^{2}-\frac{d(d-1)}{k}e-\frac{d}{k}\sum_{p\in D\cap E}\mu_{p}(E)+\sum_{[x,y,z]\in D\cap E}\mu_{[x,y,z,0]}(C)+\left(\frac{d}{k}-1\right)\#(D\cap E).$
This equality is from Theorem 4.5.
The first key obstacle for us to prove the Bounded Negativity Conjecture on $X_{d}$ comes from the term
$\sum_{p\in D\cap E}\mu_{p}(E)$
which may be comparable to $e^{2}-(d-1)e$ for $e$ sufficiently large. Moreover, this problem arises from two possible causes: (i) $D\cap E$ contains many singular points of $E$; (ii) $\mu_{p}(E)$ is large for $p\in D\cap E$.
5.1. The case #$\#(D\cap E)=de$.
We first consider the case where $D\cap E$ contains as many points as possible, i.e., #$\#(D\cap E)=de$. Then for each $p\in D\cap E$, $D$ and $E$ intersect transversely at $p$ by [6, Cor.2.6.8]. In particular, $p$ is a smooth point of $E$, and hence $\mu_{p}(E)=0$. Therefore,
$\sum_{p\in D\cap E}\mu_{p}(E)=0.$
5.2. The case where $\mu_{p}(E)$ is large for some $p\in D\cap E$.
In the extreme case, #$\#(D\cap E)=1$ and $D\cap E$ is the only singular point of $E$; moreover, $E$ is a rational cuspidal curve and $C$ is smooth. Then
$\frac{d}{k}e^{2}-\frac{d(d-1)}{k}e-\frac{d}{k}\sum_{p\in D\cap E}\mu_{p}(E)+\sum_{[x,y,z]\in D\cap E}\mu_{[x,y,z,0]}(C)+\left(\frac{d}{k}-1\right)\#(D\cap E)$
$=$ $\frac{d}{k}e^{2}-\frac{d(d-1)}{k}e-\frac{d}{k}(e-1)(e-2)+\left(\frac{d}{k}-1\right)$
$=$ $\frac{d}{k}\left((4-d)e-1\right)-1,$
which can be very negative for $e$ sufficiently large.
Hence we propose the following problem:
Problem 5.3.
Does there exist a sequence of irreducible homogeneous polynomials $g_{e_{i}}$ of degrees $e_{i}$ for a sequence of integers $e_{i}\to\infty$ such that the following conditions are satisfied?
  1. (i)
    The projective plane curve $V(g_{e_{i}})$ is a rational cuspidal curve with exactly one singular point lying on the curve $x^{d}+y^{d}+z^{d}=0$; and
  2. (ii)
    The curve $\mathscr{C}_{e_{i}}$ in $\mathbb{P}^{3}$ defined by $g_{e_{i}}(x,y,z)=0,x^{d}+y^{d}+z^{d}+w^{d}=0$ is a reducible curve with smooth irreducible components.
If the answer is positive, then taking an irreducible component $C_{e_{i}}$ of $\mathscr{C}_{e_{i}}$, we will have
$C_{e_{i}}^{2}=\frac{d}{k_{i}}\left((4-d)e_{i}-1\right)-1\to-\infty$
where $2\leq k_{i}\leq d$. This would disprove the Bounded Negativity Conjecture for $X_{d}$.
@@ 5.4. An invariant
where $V(g)$ denotes the projective plane curve defined by $g$.
Problem 5.5.
@@ 6. Puiseux series, resultants, and intersection multiplicities
In this section, we first recall basic notions concerning Puiseux series and resultants, and then use them to prove properties of intersection multiplicities of plane curve germs. Finally, we give the proof of Proposition 4.4.
@@ 6.1. Puiseux series
We follow the notations in [6, Sec.1.2, p.17]. The ring $\mathbb{C}[[u]]$ denotes the set of formal power series, and $\mathbb{C}\{u\}$ the analytic local ring of $(\mathbb{C},0)$. Denote by $\mathbb{C}((u))$ the field of fractions of $\mathbb{C}[[u]]$. A typical element of $\mathbb{C}((u))$ can be represented as
$\sum_{i\geq r}a_{i}u^{i},$
for some $r\in\mathbb{Z}$ and $a_{i}\in\mathbb{C}$.
The field of fractionary power series, denoted by $\mathbb{C}\langle\langle u\rangle\rangle$ is the direct limit of the fields $\mathbb{C}((u^{\frac{1}{n}}))$, where the morphisms in the directed system are given by
$\mathbb{C}((u^{\frac{1}{n}}))\to\mathbb{C}((u^{\frac{1}{n^{\prime}}})),\quad n|n^{\prime}\mathrm{\ and\ }\sum a_{i}u^{\frac{i}{n}}\mapsto\sum a_{i}u^{\frac{in^{\prime}/n}{n^{\prime}}}.$
Let
$s=\sum_{i\geq r}a_{i}u^{\frac{i}{n}}$
be a fractionary power series, we define the order of $s$ to be
$o_{u}(s)=\begin{cases}\infty,&\mathrm{if\ }s=0;\\ \frac{\min\{i:a_{i}\neq 0\}}{n},&\mathrm{if\ }s\neq 0.\end{cases}$
Fractionary power series $s$ with $o_{u}(s)>0$ will be called Puiseux series. After a simultaneous reduction of all the fractional exponents effectively appearing in $s$, we may assume that $n$ and $\gcd\{i:a_{i}\neq 0\}$ are coprime, and then we call $n$ the polydromy order of $s$, denoted by $\nu(s)$. We will also use the following notation: if $n=\nu(s)$ and $s$ can be written as
$s(u)=\sum a_{i}u^{\frac{i}{n}},$
then define
$s_{,\alpha}(u)=\sum e^{2\pi\sqrt{-1}\alpha i/n}a_{i}u^{\frac{i}{n}}$
for $\alpha=1,\ldots,n$. Moreover, we say that $s$ is a convergent fractionary power series if the formal power series $s(t^{n})$ has a positive convergence radius.
Suppose $f\in\mathbb{C}[[u,v]]$ is a formal power series satisfying $f(0,0)=0$. We say a Puiseux series $s\in\mathbb{C}\langle\langle u\rangle\rangle$ is a $v$-root (or simply, root) of $f$ is $f(u,s(u))=0$. Furthermore, we say that $f$ is $v$-general of order $k$, if
$f(0,v)=c_{k}v^{k}+c_{k+1}v^{k+1}+\cdots,$
for some constant $c_{k}\neq 0$.
For convenience, for $s,s^{\prime}\in\mathbb{C}\langle\langle u\rangle\rangle$, we use $s\simeq s^{\prime}$ to denote the relation
@@
for some $a\in\mathbb{C}^{*}$ and $s^{\prime\prime}$ a Puiseux series. Note that if $s\simeq s^{\prime}$, then $o_{u}(s)=o_{u}(s^{\prime})$.
Some basic results which we will implicitly use in the theory of Puiseux series are summarized as follows; we refer to [6, Chap.1] for the proof.
Proposition 6.2.
The following statements are true:
  1. (i)
    The field $\mathbb{C}\langle\langle u\rangle\rangle$ is algebraically closed.
  2. (ii)
    If $f\in\mathbb{C}\{u,v\}$ is a convergent power series and a Puiseux series $s\in\mathbb{C}\langle\langle u\rangle\rangle$ is a root of $f$, then $s$ is convergent.
  3. (iii)
    If $f\in\mathbb{C}\{u,v\}$ is $v$-general of order $b$ and a Puiseux series $s\in\mathbb{C}\langle\langle u\rangle\rangle$ is a root of $f$ such that $\nu(s)=b$, then
    $\prod_{\alpha=1}^{b}\left(v-s_{,\alpha}(u)\right)\mathrm{\ divides\ }f\mathrm{\ in\ }\mathbb{C}\{u,v\}.$
    Moreover, if, in addition, $f$ is irreducible, then
    $f=c(u,v)\prod_{\alpha=1}^{b}\left(v-s_{,\alpha}(u)\right)$
    for an invertible element $c(u,v)$ in $\mathbb{C}\{u,v\}$.
6.3. Resultants
We recall the definition and basic properties of resultants; all the results are well known. The interested reader may consult [5, Sec.1.4].
Let $K$ be a field of characteristic zero, which will be taken as $\mathbb{C}\langle\langle u\rangle\rangle$ in our discussion about intersection multiplicities. Let $f,g$ are two polynomials in $K[v]$ of the following form:
$f=a_{0}v^{d}+a_{1}v^{d-1}+\cdots+a_{d},$

### v3

@@ 3.1. Strategy
(3.1) $=m^{2}kC^{2}+m^{2}\sum_{i\neq j}C_{i}\cdot C_{j},$
where $C_{i}\cdot C_{j}$ denotes the intersection number of the curves $C_{i}$ and $C_{j}$ on $X_{d}$.
Our aim is to derive a lower bound for $C^{2}$. From formula (3.1), we are led to establish an upper bound for the sum
$\sum_{i\neq j}C_{i}\cdot C_{j},$
which will yield a lower bound for the following difference
@@
If this difference can be bounded from below as a function in $e\geq 1$ when $d$ is viewed as a constant, then the Bounded Negativity Conjecture for $X_{d}$ will follow.
In the main part of this paper, we will calculate the intersection numbers $C_{i}\cdot C_{j}$ through local computation methods. The key results are summarized in Proposition 4.4, and the lengthy proof will be presented in Section 5.
For the local computation, fix an arbitrary intersection point $P=[x_{0},y_{0},z_{0},w_{0}]$ of any pair of curves $C_{i}$ and $C_{j}$ with $i\neq j$, and we will give an explicit formula for the sum
@@
where $I_{P}(C_{i},C_{j})$ denotes the intersection multiplicity of $C_{i}$ and $C_{j}$ at the point $P$.
From $P\in X_{d}$, we have $x_{0}^{d}+y_{0}^{d}+z_{0}^{d}+w_{0}^{d}=0$, hence not all $x_{0},y_{0},z_{0}$ are zero. By the symmetry of the coordinates $x,y,z$ in the equation $X_{d}:x^{d}+y^{d}+z^{d}+w^{d}=0$, we may without loss of generality assume that $x_{0}\neq 0$. To decide where the intersection point $P=[x_{0},y_{0},z_{0},w_{0}]$ may situate on $X_{d}$, we propose the following lemma which will be proved at the end of this section.
@@
To calculate the intersection multiplicities at $P$, we introduce affine coordinates
@@
Moreover, since $x_{0}^{d}+y_{0}^{d}+z_{0}^{d}=0$ and $x_{0}\neq 0$, at least one of $y_{0},z_{0}$ is nonzero. We may without loss of generality assume that $z_{0}\neq 0$. Then
@@
Therefore, we can locally express $v$ as an analytic function of $(u,\omega)$ by applying the implicit function theorem to the equation $l(u,v)+\omega^{d}=0$. Then the local equation for $\mathscr{C}$ can be expressed as $h\left(u,v(u,\omega)\right)=0$. By analyzing this local equation, we can investigate the local branches and compute the local intersection multiplicities. This is the usual way to analyze the curve $\mathscr{C}$.
Our method is, however, first to solve $v$ from the equation $h(u,v)=0$ using Puiseux series, say $v=s(u)$, and then study the properties of the curve $\mathscr{C}$ and calculate the intersection multiplicities using the equation $l(u,s(u))+\omega^{d}=0$. To this end, we need that $h$ is $v$-general, i.e., $h(0,v)$ is not identically zero as a function in $v$.
@@
Since $h$ is transformed by an affine coordinate change from an irreducible polynomial $g$, it is an irreducible polynomial in $u,v$. If $h$ is not $v$-general, then $u$ is a factor of $h$; therefore, by irreducibility, $h$ is a constant multiple of $u$. Transforming $h$ back to $g$, we see that $g$ is a linear function yaxy-ax for some $a\in\mathbb{C}$ (up to multiplicative factor). As a result, the curve $\mathscr{C}$ is defined as
@@
If $a^{d}+1\neq 0$, then $\mathscr{C}$ is itself irreducible and hence it must be equal to $C$. We have $C^{2}=d$ in this case.
@@
Recall that we have the degree $d$ branched Galois cover $\rho:X_{d}\to\mathbb{P}^{2}$ defined by $x,y,z,w]\mapsto[x,y,z]$. The branch locus of $\rho$ is the Eermat curve $D:x^{d}+y^{d}+z^{d}=0$, and the ramification locus of $\rho$ is the inverse image $\rho^{-1}(D)$ which is defined by the equation $w=0$ on $X_{d}$.
Assume now that $Q$ is a point on at least two components of $\mathscr{C}$. Denote $q=\rho(Q)$. Then $\left(\rho^{-1}(q)\right)<d$, where #A\#A means the number of points in a set $A$. Hence $q$ lies on the branch locus $D$ and then $Q\in\rho^{-1}(D)$, so $w(Q)=0$ as desired. Furthermore, we can easily see that $\rho^{-1}(q)=\{Q\}$. For all $i\leq k$, the restriction morphism $\rho|_{C_{i}}:C_{i}\to E$ is a finite morphism and thus $\rho^{-1}(q)\cap C_{i}\neq\emptyset$. But $\rho^{-1}(q)$ consists of only one point $Q$, it follows that $Q\in C_{i}$. We are done.
@@ 4.1. Position of intersection points
By definition, $\rho^{-1}(q)=\{Q\}$ and $Q\in\mathscr{C}$. Since each $C_{i}$ is mapped onto $E$ under the branched covering map $\rho:X_{d}\to\mathbb{P}^{2}$, the intersection $\rho^{-1}(q)\cap C_{i}$ is nonempty. Hence $Q\in C_{i}$ as desired.
@@ 4.3. Intersection multiplicities
From formula (3.1) and the discussions in Section 3, to obtain a lower bound for $C^{2}$, we seek to estimate the sum
$\sum_{i\neq j}I_{P}(C_{i},C_{j}),$
where $P=[x_{0},y_{0},z_{0},0]$ with $p=[x_{0},y_{0},z_{0}]\in D\cap E$. Our key technical results are summarized as follows.
@@
    $\mu_{P}(\mathscr{C})=(d-1)i_{p}(D,E)+d\mu_{p}(E)-d+1,$
    where $i_{p}(D,E)$ denotes the intersection multiplicity of the plane curves $D$ and $E$ at the point $p$.
@@
Now we use the above results to study the Bounded Negativity Conjecture for Fermat surfaces. It is clear that we only need to consider negative curves $C$. Then the curve $\mathscr{C}$ necessarily has $k\geq 2$ irreducible components since otherwise $C^{2}$ would be nonnegative. Moreover, the number $k$ divides $d$ since the Galois action for $\rho$ is a cyclic action of order $d$. To ensure that the polynomial $h$ is $v$-general, we only need to assume $C^{2}\neq 2-d$; see the discussions in Section 3. Therefore, from Lemma 2.1 and Proposition 4.4 together with the equality
$kC^{2}=de^{2}-\sum_{i\neq j}C_{i}\cdot C_{j},$
we get the following theorem.
Theorem 4.5 (Theorem 1.1).
Let $d\geq 5$. Suppose that $C$ is a reduced and irreducible curve on $X_{d}$ with self-intersection number $C^{2}<2-d<0$. Then the following properties hold.
  1. (1)
    There exists a unique reduced and irreducible curve $E\subset\mathbb{P}^{2}$ defined by a degree $e$ polynomial $g(x,y,z)$, such that the cone in $\mathbb{P}^{3}$ determined by $g$ contains $C$.
  2. (2)
    There exists $k\geq 2$ and $k|d$ such that
@@
    In particular, when $d$ is prime, we have $k=d$ and
@@
    for some constant $\lambda>0$ which may depend on $d$; or
  2. (b)
    $\limsup_{e\to\infty}\sup_{E}\frac{\sum\limits_{p\in D\cap E}\mu_{p}(E)}{e^{2}}<1;$
where, in the above two formulae, the curve $E$ ranges over all irreducible projective plane curves of degree $e$.
@@
If (a) holds, there exists $e_{0}>d+1$ such that
@@
by [14, Cor.7.1.3], where $r_{p}$ is the number of local branches of $E$ at $p$. It follows that
@@
for some $\alpha>0$. The desired result follows from
@@ 4.7. An invariant
where $V(g)$ denotes the projective plane curve defined by $g$. We may ask the following questions which are possibly of independent interest in singularity theory of projective plane curves.
Problem 4.8.
@@ 5. Puiseux series, resultants, and intersection multiplicities
In this section, we give the proof of Proposition 4.4, by using Puiseux series and resultants. Since these techniques have become standard tools in algebraic curve theory, with well-established references available. For conciseness, we just introduce simplified notations adapted to our context and refer to standard books for comprehensive definitions and properties.
@@ 5.1. Puiseux series
For standard definitions, notations and properties of Puiseux series, we refer to [6, Chap.1]. For convenience, for a non-zero Puiseux series
$s(u)=\sum_{i=1}^{\infty}a_{i}u^{\frac{i}{n}},$
with polydromy order $\nu(s)=n$, we set
$s_{,\alpha}(u)=\sum_{i=1}^{\infty}\exp\left(\frac{2\pi\sqrt{-1}\alpha i}{n}\right)a_{i}u^{\frac{i}{n}}$
for $\alpha=1,\ldots,n$; this notation will be used to simplify the expression of decomposing an irreducible polynomial into a product of Puiseux series. Moreover, for $s,s^{\prime}\in\mathbb{C}\langle\hskip-1.5pt\langle u\rangle\hskip-1.5pt\rangle$, denote $s\simeq s^{\prime}$ if
@@
for some $a\in\mathbb{C}^{*}$ and $s^{\prime\prime}$ a Puiseux series; we will only use the property that $s\simeq s^{\prime}$ implies $o_{u}(s)=o_{u}(s^{\prime})$ (here, $o_{u}(s)$ denotes the order of $s$ in $u$ [6, Sec.1.2]).
5.2. Resultants, discriminants, and intersection multiplicities
The references abound. For the definition and basic properties, the interested reader may consult [5, Sec.1.4], [8, Chap.12],[6, Chap.2], [11, Chap.I, Sec.3], and [14, Chap.1]. The notations used in different references varying a little, we will fix notations that are suited for our purpose.
Suppose $F,G$ are two reduced curve germs in $(\mathbb{C}^{2},0)$, defined by
F:f(u,v)=0,F:\ f(u,v)=0,